OpenStudy (anonymous):

4x+3y=55 an 5x-3y=35 substitution method???

OpenStudy (anonymous):

4x+3y=55---(a) 5x-3y=35---(b) from equation (b) y = (5x-35)/3 substitute this expression of y in equation (a) 4x+3(5x-35)/3 = 55 4x+(5x-35)=55 9x-35=55 9x=90 x=10 put x=10 in (a) 4(10)+3y=55 3y=55-40=15 y=15/3 =5 y=5 x=10 y=5

OpenStudy (anonymous):

find y for 2x+2y=2 and x+2y=1 ?????

OpenStudy (anonymous):

2x+2y=2---(a) x+2y=1 ---(b) multiply eq (b) with 2 and subtract from eqn (a) 2x+2y=2---(a) (-) 2x+4y=2 ---------------- 0-2y = 0 y=0